A fishing boat traveled 3 hours against a 6km/hr current. The Return trip took only 2 hours. Find the speed of the boat in still water.
Answer by Electrified_Levi(103)

Hi, Hope I can help
A fishing boat traveled 3 hours against a 6km/hr current. The Return trip took only 2 hours. Find the speed of the boat in still water.
.
Distance=(rate)(time)
It can also be put like this.
(rate)(time)=(rate)(time)
We already know the time 3 hours, and 2 hours
We have to find the speed of the boat.
.
speed of the boat in still water = "x"
speed of the boat against current = (x - 6)(current slows the boat's speed)
speed of the boat with the current = (x + 6)(current speeds up the boat's speed)
.
Now we can solve "x"
We know both the speed(rate) and the time.
.
(rate)(time)=(rate)(time)
.
(x - 6)(3)= (x + 6)(2)
.
3(x - 6)= 2(x + 6)
.
We will use the distributive property
.
3x - 18 = 2x + 12
.
We will subtract 2x from both sides
.
x - 18 = 0 + 12
.
Now add 18 to both sides
.
x - 0 = 18 + 12
.
x = 30
.
The fishing boat went 30 km./hr in still water
.
3(30 - 6)= 2(30 + 6)
3(24)=2(36)
(72)=(72)
.
It went 72 miles each way, 144 miles all together
.
Hope I helped, Levi